OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (−2, 6) and has a slope of −1? x + y = −4 x + y = 4 x − y = −4 x − y = 4

OpenStudy (anonymous):

y=-1x+b 6=-1(-2)+b b=4

OpenStudy (anonymous):

so the equation is y=-1x+4

OpenStudy (anonymous):

x+y = 4

OpenStudy (anonymous):

so is B

OpenStudy (anonymous):

yepp